OpenStudy (vera_ewing):

What is the pH of a 0.025 M solution of NaOH? A. 2.50 B. 1.60 C. 12.40 D. Cannot determine without the Kb of NaOH

OpenStudy (vera_ewing):

@taramgrant0543664

OpenStudy (photon336):

0.025 = [OH] [OH][H+] = 1x10^-14 (1x10^-14)/(0.025) = 4x10^-13 -log(4x10^-13) =12.4

OpenStudy (taramgrant0543664):

For this one you solve the same way that you normally would and add an extra step at the end since it's OH To start we use the formula: pOH=-log[OH] =-log(0.035) =1.602 Since it's asking for pH and not pOH we just do a simple subtraction: pH=14-pOH =14-1.602 =12.398

OpenStudy (vera_ewing):

Thank you! That makes so much more sense now! :)

OpenStudy (taramgrant0543664):

No problem!!

OpenStudy (photon336):

Explanation: You're given the concentration of OH- which you need to use to find the pH pH = -log(H) which p means take the -log of something. At 25 degrees Celsius you know that the [H][OH] = 10^-14 so if you know the concentration of one you know the concentration of the other.